# An Asymptotically Optimal Coordinate Descent Algorithm for Learning Bayesian Networks from Gaussian Models

### v1

@@ 4.2. Statistical Consistency Guarantees
Here, entries of $\tilde{\Gamma}$ lie between $\hat{\Gamma}$ and $\hat{\Gamma}^{\mathrm{opt}}$, and entries of $\bar{\Gamma}$ lie between $\hat{\Gamma}^{\mathrm{opt}}$ and $\Gamma^{\star}$. Some algebra then gives:
$1/2(\hat{\Gamma}-\hat{\Gamma}^{\mathrm{opt}})^{\mathrm{\scriptscriptstyle T}}\nabla^{2}{\ell}(\tilde{\Gamma})(\hat{\Gamma}-\hat{\Gamma}^{\mathrm{opt}})\leq&[\ell(\hat{\Gamma})-\ell(\hat{\Gamma}^{\mathrm{opt}})]+\|\nabla\ell(\Gamma^{\star})\|_{\ell_{2}}\|\hat{\Gamma}-\hat{\Gamma}^{\mathrm{opt}}\|_{\ell_{2}}\\ &+\|\hat{\Gamma}-\hat{\Gamma}^{\mathrm{opt}}\|_{\ell_{2}}\|\hat{\Gamma}^{\mathrm{opt}}-\Gamma^{\star}\|_{\ell_{2}}\kappa_{\text{max}}(\nabla^{2}\ell(\bar{\Gamma})).\end{aligned}$
By the convexity of $\ell(\cdot)$, for any $\Gamma$, $\nabla^{2}\ell(\Gamma)\succeq\hat{\Sigma}\otimes I$. Thus appealing to Lemma 10, with probability greater than $1-\mathcal{O}(1/n)$, $\sigma_{\text{min}}(\nabla^{2}\ell(\Gamma))\geq\underline{\kappa}^{2}/2$. Letting $\tau:=4(\|\hat{\Gamma}^{\mathrm{opt}}-\Gamma^{\star}\|_{\ell_{2}}\kappa_{\text{max}}(\nabla^{2}\ell(\bar{\Gamma}))+\|\nabla\ell(\Gamma^{\star})\|_{\ell_{2}})/\underline{\kappa}^{2}$, with probability greater than $1-\mathcal{O}(1/n)$: $\|\hat{\Gamma}-\hat{\Gamma}^{\mathrm{opt}}\|_{\ell_{2}}^{2}\leq 4\underline{\kappa}^{-2}\ell(\hat{\Gamma})-\ell(\hat{\Gamma}^{\mathrm{opt}})+4\tau\underline{\kappa}^{-2}\|\hat{\Gamma}-\hat{\Gamma}^{\mathrm{opt}}\|_{\ell_{2}}\tau$. Note that for non-negative $Z,W,\Pi$, the inequality $Z^{2}\leq\Pi{Z}+W$ implies $Z\leq(\Pi+\sqrt{\Pi^{2}+4W})/2$. Using this fact, in conjunction with the previous bound, we obtain with probability greater than $1-\mathcal{O}(1/n)$ the bound $\|\hat{\Gamma}-\hat{\Gamma}^{\mathrm{opt}}\|_{\ell_{2}}\leq\frac{\tau}{2}+\frac{1}{2}(\tau^{2}+16\underline{\kappa}^{-2}[\ell(\hat{\Gamma})-\ell(\hat{\Gamma}^{\mathrm{opt}})])^{1/2}$. We next bound $\tau$. From Corollary 13, we have control over the term $\|\hat{\Gamma}^{\mathrm{opt}}-\Gamma^{\star}\|_{\ell_{2}}$ in $\tau$. It remains to control $\sigma_{\text{max}}(\nabla^{2}\ell(\bar{\Gamma}))$ and $\|\nabla\ell(\Gamma^{\star})\|_{\ell_{2}}$. Let $\Gamma\in\mathbb{R}^{m^{2}}$. Suppose that for every $j\in T$, $\Gamma_{j}\geq\nu$. Then, some calculations yield the bound $\nabla^{2}\ell(\Gamma)\preceq\hat{\Sigma}\otimes I+\frac{2}{\nu^{2}}{I}_{m^{2}}=\hat{\Sigma}\otimes I+\frac{2}{\nu^{2}}{I}_{m^{2}}$. We have that for every $j\in T$, $\hat{\Gamma}^{\mathrm{opt}}_{j}\geq\Gamma^{\star}_{j}-\|\hat{\Gamma}^{\mathrm{opt}}-\Gamma^{\star}\|_{\ell_{2}}$. From Corollary 13, Assumption 5, and that $\lambda\sqrt{s^{\star}}\leq 1$, we then have $\hat{\Gamma}^{\mathrm{opt}}_{j}\geq\Gamma^{\star}_{j}/{2}\geq 1/2(\Omega^{\star}_{j})^{-1/2}$. Since the entries of $\bar{\Gamma}$ are between those of $\Gamma^{\star}$ and $\hat{\Gamma}^{\mathrm{opt}}$ and by Lemma 10, $\sigma_{\text{max}}(\nabla^{2}\ell(\bar{\Gamma}))\leq\sigma_{\text{max}}(\hat{\Sigma})+8\min_{j}\Omega^{\star}_{j}=\mathcal{O}(1)$. To control $\nabla\ell(\Gamma^{\star})$, we first note that $\mathbb{E}[\nabla\ell(\Gamma^{\star})]=0$. Therefore, $\|\nabla\ell(\Gamma^{\star})\|_{\ell_{2}}=\|\nabla\ell(\Gamma^{\star})-\mathbb{E}[\nabla\ell(\Gamma^{\star})]\|_{\ell_{2}}$. Since $\nabla\ell(\Gamma^{\star})-\mathbb{E}[\nabla\ell(\Gamma^{\star})]=((\hat{\Sigma}-\Sigma^{\star})\otimes I)\Gamma^{\star}$, letting $K^{\star}=({\Sigma^{\star}})^{-1}$ we get $\|\nabla\ell(\Gamma^{\star})-\mathbb{E}[\nabla\ell(\Gamma^{\star})]\|_{\ell_{2}}^{2}=\mathrm{tr}((\hat{\Sigma}_{n}-\Sigma^{\star})(\hat{\Sigma}_{n}-\Sigma^{\star})^{\mathrm{\scriptscriptstyle T}}K^{\star})\leq\|\hat{\Sigma}-\Sigma^{\star}\|_{2}^{2}\|K^{\star}\|_{\star}\leq m\|\hat{\Sigma}-\Sigma^{\star}\|_{2}^{2}\|K^{\star}\|_{2}\leq\mathcal{O}({m^{2}\log(n)}/{n})$. Thus, $\|\nabla\ell(\Gamma^{\star})-\mathbb{E}[\nabla\ell(\Gamma^{\star})]\|_{\ell_{2}}\leq\mathcal{O}({m}\sqrt{\log{n}}/{\sqrt{n}})$. Upper bounding $\tau$ and then ultimately using that to upper-bound $\|\hat{\Gamma}-\hat{\Gamma}^{\mathrm{opt}}\|_{\ell_{2}}$, we conclude that $\|\hat{\Gamma}-\hat{\Gamma}^{\mathrm{opt}}\|_{\ell_{2}}^{2}\leq\mathcal{O}(\sqrt{d_{\mathrm{max}}^{2}m^{4}\log{m}/n})$. Combining this bound with Proposition 12, we get the first result of the theorem. The second result follows straightforwardly from triangle inequality: $\|\hat{\Gamma}-\Gamma^{\star}\|_{F}^{2}\leq 2\|\hat{\Gamma}-\hat{\Gamma}^{\mathrm{opt}}\|_{F}^{2}+2\|\hat{\Gamma}^{\mathrm{opt}}-\Gamma^{\star}\|_{F}^{2}\leq\mathcal{O}(\sqrt{d_{\mathrm{max}}^{2}m^{4}\log{m}/n})$. ∎
The result of Theorem 7 guarantees that the estimate from our coordinate descent procedure is close to the optimal solution of (3), and that it accurately estimates certain reordering of the population model. For accurately estimating the edges of the population Markov equivalence class $\mathrm{MEC}(\mathcal{G}^{\star})$, we need the faithfulness condition and a strictly stronger version of the beta-min condition[23], dubbed the strong beta-min condition.
Assumption 7.
(Faithfulness)The DAG $\mathcal{G}^{\star}$ is faithful with respect to the data generating distribution $\mathcal{P}^{\star}$, that is, every conditional independence relationship entailed in $\mathcal{P}^{\star}$ is encoded $\mathcal{G}^{\star}$.
Assumption 8.
(Strong beta-min condition) There exist constant $0<\eta_{0}^{2}<1/s^{\star}$, such that for any $\pi\in\Pi$, the matrix $\tilde{B}^{\star}(\pi)$ has all of its nonzero coordinates $(k,j)$ satisfy $|\tilde{B}^{\star}_{kj}(\pi)|>\sqrt{s^{\star}\log{m}/n}/\eta_{0}$.
Theorem 14.
Suppose $\lambda^{2}\asymp s^{\star}\log{m}/n$, the sample size satisfies $n/\log(n)\geq\mathcal{O}(m^{2}\log{m})$, and assumptions of Theorem 11 hold, with Assumption 4 replaced by Assumption 8. Then, with probability greater than $1-2\alpha_{0}$, there exists a member of the population Markov equivalence class with associated parameter $\Gamma^{\star}_{\mathrm{mec}}$ such that $\|\hat{\Gamma}-\Gamma^{\star}_{\mathrm{mec}}\|_{F}^{2}\leq\mathcal{O}(\sqrt{d_{\text{max}}^{2}m^{4}\log{m}/n})$.
Appealing to Remark 3.2 of van de Geer and Bühlmann [23], under assumptions of Theorem 11, as well as Assumption 8, the graph encoded by any optimal connectivity matrix $\hat{B}^{\text{opt}}$ of this optimization problem encodes, with probability $1-\alpha_{0}$, a member of the Markov equivalence class of the population directed acyclic graph. Let $(B_{\mathrm{mec}}^{\star},\Omega_{\mathrm{mec}}^{\star})$ be the associated connectivity matrix and noise matrix of this population model. Furthermore, define $\Gamma_{\mathrm{mec}}^{\star}=(I-B_{\mathrm{mec}}^{\star}){\Omega_{\mathrm{mec}}^{\star}}^{-1/2}$. The proof of the theorem relies on the following lemma in [25].
Lemma 15 (Lemma 7 of [25]).
Under the conditions of Theorem 14, we have with probability greater than $1-2\alpha_{0}$, $\|\hat{\Gamma}^{\mathrm{opt}}-\Gamma^{\star}_{\mathrm{mec}}\|_{F}^{2}=\mathcal{O}({m^{2}}/{n}).$
Proof of Theorem 14.
First, by Lemma 15, with probability greater than $1-2\alpha_{0}$, $\|\hat{\Gamma}-\Gamma^{\star}_{\mathrm{mec}}\|_{F}^{2}\leq 2\|\hat{\Gamma}-\hat{\Gamma}^{\mathrm{opt}}\|_{F}^{2}+2\|\hat{\Gamma}^{\mathrm{opt}}-\Gamma^{\star}_{\mathrm{mec}}\|_{F}^{2}\leq\text{GAP}+\mathcal{O}({m^{2}}/{n})$. Since the GAP is on the order $\mathcal{O}(\sqrt{d_{\mathrm{max}}^{2}m^{4}\log{m}/n})$, we get $\|\hat{\Gamma}-\Gamma^{\star}_{\mathrm{mec}}\|_{F}^{2}\leq\mathcal{O}(\sqrt{d_{\mathrm{max}}^{2}m^{4}\log{m}/n})$. ∎
We remark that without the faithfulness condition (see Assumption 7), we can guarantee that the estimate from our coordinate descent procedure is close to a member of what is known as the minimal-edge I-MAP. The minimal-edge I-MAP is the sparsest set of directed acyclic graphs that induce a structural equation model compatible with the true data distribution. Under faithfulness, the minimal-edge I-MAP coincides with the population Markov equivalence class [23].
5. Experiments
@@
In this section, we illustrate the utility of our method on synthetic and real data and compare its performance with competing methods. We dub our method CD-$\ell_{0}$ as it is a coordinate descent method using $\ell_{0}$ penalized loss function. The competing methods we compare against include Greedy equivalence search (GES) [5], Greedy Sparsest Permutation (GSP) [19], and the mixed-integer convex program (MICODAG) [25]. We also compare our method with other coordinate descent algorithms (CCDr-MCP) [1, 2, 9], which use a minimax concave penalty instead of $\ell_{0}$ norm and are implemented as an R package sparsebn. All experiments are performed with a MacBook Air (M2 chip) with 8GB of RAM and a 256GB SSD, using Gurobi 10.0.0 as the optimization solver.
@@
As the input super-structure $E_{\text{super}}$, we supply an estimated moral graph, computed using the graphical lasso procedure [8]. To make our comparisons fair, we appropriately modify the competing methods so that $E_{\text{super}}$ can also be supplied as input. Note that we count the number of support after each update in Algorithm 1. Converting the graph into a string key at each iteration is inefficient. Therefore, in the implementation, we count the support only after each full loop, setting the threshold to $C$ instead of $Cm^{2}$. Throughout this paper, $C$ is set to 5.
@@
The time limit for the integer programming method MICODAG is set to $50m$. If the algorithm does not terminate within the time limit, we report the solution time (in seconds) and the achieved relative optimality gap, computed as RGAP=(upper boundlower bound)/lower $\mathrm{RGAP}=(\text{upper bound}-\text{lower bound})/\text{lower bound}$. Here, the upper bound and lower bound refer to the objective value associated with the best feasible solution and best lower bound, obtained respectively by MICODAG. A zero value for $\mathrm{RGAP}$ indicates that an optimal solution has been found.
Unless stated otherwise, we use the Bayesian information criterion (BIC) to choose the parameter $\lambda$. In our context, the BIC score is given by $-2n\sum_{i=1}^{m}\log(\hat{\Gamma}_{ii})+n\mathrm{tr}(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})+k\log(n)$, where $k$ is the number of nonzero entries in the estimated parameter $\hat{\Gamma}$. From theoretical guarantees in [25], $\lambda^{2}$ should be on the order ${\log(m)/n}$. Hence, we choose $\lambda$ with the smallest BIC score among $\lambda^{2}=c\log{m}/n$, for $c=1,2,\ldots,15$.

### v3

@@ 4.2 Optimality guarantees
[…]
Figure 1: Left: scenario for Setting I, middle: scenario for setting II.1, and right: scenario for setting II.2; solid directed edges represent directed edges that are assumed to be in the estimate $\hat{E}$, crossed out solid directed edges represent directed edges that are assumed to be excluded in the estimate $\hat{E}$, crossed out solid undirected edges indicate that the corresponding nodes are not connected in $\hat{E}$, and crossed out dashed directed edge indicates that the edge is not present in $\hat{E}$ as adding it would create a cycle.
Part 2) When the sample size $n$ is finite, the entries of the nonzero value of $\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}$ that would be zero in the ideal DAG structure are instead small but nonzero, due to estimation error in the sample covariance matrix. We use the matrix $\Delta$ to capture these small perturbations. Then, using the proof of Theorem 7 part i), we can immediately conclude that the matrix $\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}_{n}$ can be decomposed as the sum $N+\Delta$. Here, the off-diagonal entries of $N$ specify a DAG, with ones on the diagonal and under the assumption on $n$, with probability greater than $1-\mathcal{O}(1/n)$, $\|\Delta\|_{\infty}\leq 4\bar{\kappa}(1+d_{\text{max}})\lambda$ with zeros on the diagonal of $\Delta$ (here, $\|\cdot\|_{\infty}$ denotes the maximum entry of the input matrix in absolute value). Consequently, $\|\Delta\|_{2}\leq 4m\bar{\kappa}(1+d_{\text{max}})\lambda$. Furthermore, by Lemma 10, we get $\sigma_{\text{min}}(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}_{n})\geq\sigma_{\text{min}}(\hat{\Gamma})^{2}\sigma_{\text{min}}(\hat{\Sigma})\geq\underline{\kappa}^{4}/4$. The reverse triangle inequality yields $\sigma_{\text{min}}(N)\geq\underline{\kappa}^{4}/4-4m\bar{\kappa}^{2}(1+d_{\text{max}})\lambda$. Consider any matrix $\bar{N}$ with $|\bar{N}_{ij}-N_{ij}|\leq|\Delta_{ij}|$. Using the reverse triangle inequality again, we get $\sigma_{\text{min}}(\bar{N})\geq\underline{\kappa}^{4}-8m\bar{\kappa}^{2}(1+d_{\text{max}})\lambda$ with probability greater than $1-\mathcal{O}(1/n)$. By the assumption that the sample size satisfies $n/\log(n)\geq\mathcal{O}(m^{2}\log{m})$, $\bar{N}$ is invertible, and so we can use the first-order Taylor series expansion to obtain $\log\det(N+\Delta)=-\log\det(N)-\mathrm{tr}(\bar{N}^{-1}\Delta)$. Since $\log\det(N)=0$, we obtain the bound $\log\det(N+\Delta)\leq-\mathrm{tr}(\bar{N}^{-1}\Delta)\leq\|\bar{N}^{-1}\|_{2}\|\Delta\|_{\star}$ with $\|\cdot\|_{\star}$ denoting the nuclear norm. Thus, $\log\det(N+\Delta)\leq\|\bar{N}^{-1}\|_{2}\|\Delta\|_{\star}\leq\frac{m}{\sigma_{\text{min}}(\bar{N})}\|\Delta\|_{2}\leq\frac{4m^{2}\bar{\kappa}^{2}(1+d_{\text{max}})\lambda}{\underline{\kappa}^{4}/4-8m\bar{\kappa}^{2}(1+d_{\text{max}})\lambda}$. As $\lambda^{2}=\mathcal{O}(\log{m}/n)$, by the assumption on the sample size, $f(\hat{\Gamma})-f(\hat{\Gamma}^{\mathrm{opt}})\leq\mathcal{O}(\sqrt{d_{\mathrm{max}}^{2}m^{4}\log{m}/n})$.
4.3 On the Optimization Landscape of Problem (3)
The optimization problem (3) is highly non-convex, due to the $\ell_{0}$ penalty as well as the DAG constraint. As a result, its landscape may contain many local minimizers. The following corollary to Theorem 7 analyzes these local minimizers.
Corollary 11
Consider the setup in Theorem 7. Let $\tilde{\Gamma}$ be any local minimizer to (3). Then, $f(\tilde{\Gamma})-f(\hat{\Gamma}^{\mathrm{opt}})\to_{P}0$ as $n\rightarrow\infty$. Furthermore, if $n/\log(n)\geq\mathcal{O}(m^{2}\log{m})$, with probability greater than $1-1/\mathcal{O}(n)$, we have $0\leq f(\tilde{\Gamma})-f(\hat{\Gamma}^{\mathrm{opt}})\leq\mathcal{O}(\sqrt{d_{\text{max}}^{2}m^{4}\log{m}/n})$.
This corollary follows from Theorem 7 and the fact that coordinate-wise optimality is weaker than local optimality; therefore, any local minima of this problem satisfies the same result. This result implies that when the sample size is large, all local minima achieve objective values close to a global minimum; moreover, in the infinite sample regime, all local minimizers are global minimizers. Given the non-convexity of our objective function, this result offers significant insight into the optimization landscape of the problem.
5 Selecting a Suitable Ordering for Coordinate Descent Updates
A significant challenge for nonconvex optimization problems is that selecting a good starting point and/or an effective coordinate update order is crucial for achieving good solutions. For example, in our setting, the performance of Algorithm 1 can be sensitive to the input ordering $O$ for coordinate descent updates. To address this challenge, we propose estimating a topological ordering of the underlying DAG $\mathcal{G}^{\star}$ and using it as the input ordering $O$. A valid topological ordering of $\mathcal{G}^{\star}$ is a permutation $(a_{1},a_{2},\dots,a_{m})$ of $(1,2,\dots,m)$ such that for every edge $i\to j$ in $\mathcal{G}^{\star}$, the index $a_{i}$ precedes $a_{j}$ in the ordering.
To estimate a good topological ordering, we use an algorithm proposed by Chen et al. (2019). This procedure, outlined in Algorithm 3, iteratively selects the node with the smallest conditional variance given the variables selected in the previous step. When the variances of the noise (i.e., variance of the coordinates of $\epsilon$ in (1)) are identical, they show that with high probability, their algorithm correctly recovers a valid topological ordering of $\mathcal{G}^{\star}$.
We improve upon the theoretical results of Chen et al. (2019) and show that we can still use Algorithm 3 to recover a valid topological ordering, even when the noise variances are not identical. Our result assumes that the noise variances are not too far apart—that is, the noise distribution is nearly homoscedastic.
Algorithm 3 Topological Ordering (Chen et al., 2019)
1:Input: $\hat{\Sigma}\in\mathbb{R}^{p\times p}$ (estimated) covariance of $X$
[…]
3:for $t=1,\ldots,m$ do
4:  $j^{*}\leftarrow\arg\min\limits_{j\in V\setminus C}\hat{\Sigma}_{j,j}-\hat{\Sigma}_{j,C}\hat{\Sigma}_{C,C}^{-1}\hat{\Sigma}_{C,j}=\frac{1}{\left\{(\hat{\Sigma}_{C\cup\{j\},C\cup\{j\}})^{-1}\right\}_{j,j}}$
5:  Append $j^{*}$ to $C$
6:end for
7:Output: $C$
Assumption 3 (Near homoscedastic noise)
Define $\zeta:=\min_{(j,k)\in E}(B^{\star}_{jk})^{2}$. Let $\Omega^{\star}_{\max}:=\max_{j}\Omega^{\star}_{j,j}$ and $\Omega^{\star}_{\min}:=\min_{j}\Omega^{\star}_{j,j}$ denote the maximum and minimum error variance, respectively. We assume $\Omega^{\star}_{\max}<\Omega^{\star}_{\min}(1+\zeta)$; that is, the maximum and minimum error variances are not too far apart.
Theorem 12
Let $X$ be variables defined in (1). Let $\lambda_{\min}>0$ be the smallest eigenvalue of the population covariance matrix $\Sigma^{\star}$. Suppose Assumption 3 holds. If
$n>\;$ $128m^{2}\log\left(\frac{2m^{2}+2m}{\epsilon}\right)\left(1+4\frac{\Omega^{\star}_{\max}}{\Omega^{\star}_{\min}}\right)^{2}\left(\max_{j\in V}\Sigma^{\star}_{j,j}\right)^{2}$
$\times\left(\frac{1}{\lambda_{\min}}+\frac{2\Omega^{\star}_{\max}\Omega^{\star}_{\min}(1+\zeta)}{(\Omega^{\star}_{\min}(1+\zeta)-\Omega^{\star}_{\max})\lambda^{2}_{\min}}\right)^{2},$
then Algorithm 3 recovers a valid topological ordering of $\mathcal{G}^{\star}$ with probability at least $1-\epsilon$.
The proof of Theorem 12, presented in Appendix B, adapts the argument from Chen et al. (2019). Theorem 12 shows that, under Assumption 3 and with a sufficiently large sample size, Algorithm 3 recovers the correct topological ordering with high probability. Unlike the result in Chen et al. (2019), which assumes equal noise variances (homoscedasticity), our analysis allows for heteroscedastic noise, provided the variances are not too dissimilar, as formalized in Assumption 3. This generalization is important in practice, where the assumption of exactly equal noise variances can be unrealistic.
Remark 13 (Statistical consistency.)
Suppose a valid topological ordering of the $m$ variables is given. Then, under such ordering, the true weighted adjacency matrix $\Gamma^{\star}$ of the DAG is strictly upper‐triangular, and the population covariance satisfies $\Sigma^{\star}={\Gamma^{\star}}^{{\mathrm{\scriptscriptstyle-T}}}\,{\Gamma^{\star}}^{-1}$. This implies that $\Gamma^{\star}={L^{\star}}^{\mathrm{\scriptscriptstyle-T}}$, where $L^{\star}$ is the Cholesky factor of $\Sigma^{\star}$, i.e., $\Sigma^{\star}=L^{\star}{L^{\star}}^{\mathrm{\scriptscriptstyle T}}$. Thus, the problem of finding the DAG reduces to the known sparse Cholesky factor estimation problem. In fact, consistency (in Frobenius or operator norm) of the sparse Cholesky factor follows directly from existing results on regularized Cholesky‐factor estimation; see, for example, Theorem 10 of Lam and Fan (2009).
6 Synthetic and Real Experiments
@@
In this section, we illustrate the utility of our method in Algorithm 1 on synthetic and real data and compare its performance with competing methods. We dub our method CD-$\ell_{0}$ as it is a coordinate descent method using $\ell_{0}$ penalized loss function. The competing methods we compare against include greedy equivalence search (GES) (Chickering, 2002), NOTEARS (Zheng et al., 2018), and the mixed-integer convex program (MICODAG) (Xu et al., 2024). We also compare our method with other coordinate descent algorithms (CCDr-MCP) (Aragam and Zhou, 2015; Aragam et al., 2019; Fu and Zhou, 2013), which use a minimax concave penalty instead of $\ell_{0}$ norm and are implemented as an R package sparsebn. To obtain an input ordering $O$ for CD-$\ell_{0}$, we use the algorithm of Chen et al. (2019) (presented in Algorithm 3). Consequently, we include in our comparisons the method of Chen et al. (2019), which performs statistical testing to infer a network structure using this estimated ordering. All experiments were conducted on an Intel Xeon Gold 6230R CPU with eight cores and 8 GB of memory, using Gurobi 10.0.1 as the optimization solver.
@@
As the input super-structure $E_{\text{super}}$, we supply an estimated moral graph, computed using the graphical lasso procedure (Friedman et al., 2008). In all experiments, we set the graphical lasso penalty to $0.01$ (using a larger value for some datasets to avoid ill-conditioning) and then threshold the estimated precision matrix by retaining only those entries with absolute value $\geq 0.1$. To make our comparisons fair, we appropriately modify the competing methods so that $E_{\text{super}}$ can also be supplied as input. Note that we count the number of support after each update in Algorithm 1. Converting the graph into a string key at each iteration is inefficient. Therefore, in the implementation, we count the support only after each full loop, setting the threshold to $C$ instead of $Cm^{2}$. Throughout this paper, $C$ is set to 5. We experimented with different initializations of $\Gamma^{\mathrm{init}}$ beyond the default value and observed similar performance across the various choices. As a result, throughout this section, we use the default $\Gamma^{\mathrm{init}}=\text{identity}$. We emphasize again that our theoretical convergence and optimality results in Section 4 hold for any initialization.
@@
The time limit for the integer programming method MICODAG is set to $50m$. If the algorithm does not terminate within the time limit, we report the solution time (in seconds) and the achieved relative optimality gap, computed as $\mathrm{RGAP}=(\text{upper bound}-\text{lower bound})/\text{lower bound}$. Here, the upper bound and lower bound refer to the objective value associated with the best feasible solution and best lower bound, obtained respectively by MICODAG. A zero value for $\mathrm{RGAP}$ indicates that an optimal solution has been found.
For each method, we report results under oracle tuning, i.e., selecting the parameter that yields the best performance. This provides a benchmark of the best-case potential of the methods.